The cost in dollars to ship a package “Next Day” of one pound using Federal Express can be calculated using the linear equation:
y + 2.085x + 15.08
where x is the weight (pounds) of the package. Draw the graphic of the linear equation of the cost of shipping if the weight is 6.5 pounds

Answer by Boreal(15235)

y=2.085x + 15.08
edited
y=2.085(6.5)+15.08=13.55+15.08=$28.63